Operations Report – May 31st

SOL: 31May2018

Name of person filing report: Shannon Rupert

Non-nominal systems: Only me! Only me!

Notes on non-nominal systems: This morning I took on the power system again. I changed two settings on the AC controls to what I believed meant the solar took precedence over the generator for delivery of power to the generator. I had been working on this system without any improvement for so long that I didn’t give it another thought except that I would have to keep running the generator as the SOC fell during the day. But it never fell! We are back on solar power. I’m not declaring complete victory yet, but I am hopeful. In addition, the contractor will be here on Saturday to go over the system with me.

Generator (hours run): Turned off at 8 am, will turn back on at 8 pm

Solar—100 % SOC

Diesel – not checked %

Propane – not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –Not checked

Water (trailer) – 0 gallons.

Water (static) – not checked gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: yes. I intended to flush it again and have no one use it until the tv crew arrives. But people used it anyway. So I have to do it again. Sigh.

ATV’s Used: Yes

Oil Added? Yes

ATV Fuel Used: n/a

# Hours the ATVs were Used today: 1

Notes on ATVs: We just used them to block the driveway today so people did not drive up to the Hab.

Deimos rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? No

General notes and comments: AC to GreenHab completed. AC in Hab not yet installed.

Today was the first full day of the URC and my judging duties kept me from working on the things I had hoped to do. They will have to wait, too tired tonight to do anything more.

Summary of internet: Nothing to report

Summary of suits and radios: Suits in RAM. Six remaining new radios put into service.

Summary of Hab operations: Chaos reigns.

Summary of GreenHab operations: Water was hooked up to the swamp cooler so we now have AC in the greenhouse. URC operations on lower deck.

Summary of ScienceDome operations: URC operations happened today. We discovered that the fridge in the lab is filled with moldy petri dishes. I will clean it out in the morning before we start the URC events,

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time